OpenStudy (cherie_magee):

0.01x - 0.3y = 1 y = 0.1x - 2 Which of the following systems is equivalent to the given system? x - 30y = 100 and 10y = x - 20 x - 3y = 10 and 10y = x - 20 10x - 30y = 10 and 10y = x - 20

hartnn (hartnn):

multiply each term of 1st equation by 100, what u get ?

OpenStudy (cherie_magee):

idk @hartnn

hartnn (hartnn):

what is 0.01 * 100 = ... ?

OpenStudy (cherie_magee):

1 @hartnn

hartnn (hartnn):

correct! -0.3 * 100 = ... ?

OpenStudy (cherie_magee):

-30 @hartnn

hartnn (hartnn):

good! so the 1st equation becomes x-30y=100 right ? multiply 2nd equation by 10, what u get ?

hartnn (hartnn):

or since you have choices, you'd be able to figure out the correct one now ;)